A capacitor consists of two closely spaced metal conductors of large area, separated by a thin insulating foil. It has an electrical capacity of 3000.0 μF and is charged to a potential difference of 60.0 V. Calculate the amount of energy stored in the capacitor.

Tries 0/99

Calculate the charge on this capacitor when the electrical energy stored in the capacitor is 5.40 J.

Tries 0/99

If the two plates of the capacitor have their separation increased by a factor of 5 while the charge on the plates remains constant, by what factor is the energy stored in the capacitor increased?
Answer #1

Solution-

a.) energy = (1/2) C V2 = (1/2) * 3000.0 *10^-6 * 60^2

= 5.4 Joules
b.) energy = (1/2) Q2 / C

Q = ( 2 E C )^1/2
    = ( 2 * 5.40 * 3000 x10^-6 )^1/2

= 0.18 coulombs
c.) the capacitance decreased by a factor of 5 and therefore the energy increases by a factor of 5